# Shift-cyclicity in Analytic Function Spaces

### v1

@@ 3.3. On the unit ball $\mathbb{B}_{d}$
Fix $d\in\mathbb{N}$ and each $\alpha\in\mathbb{Z}_{+}^{d}$, we write
@@
For each $t\in\mathbb{R}$, we define the Dirichlet-type spaces on $\mathbb{B}_{d}$ as
@@
While we did not introduce these spaces formally, it is known that $\mathcal{D}_{t}(\mathbb{B}_{d})$ corresponds to the Bergman, Hardy, Drury–Arveson and Dirichlet spaces on $\mathbb{B}_{d}$ for $t=-1,0,d-1$ and $d$ respectively. Moreover, $\mathcal{D}_{t}(\mathbb{B}_{d})$ is a subalgebra of functions in $\operatorname{Hol}(\mathbb{B}_{d})$ that are continuous up to $\overline{\mathbb{B}_{d}}$ whenever $t>d$. For more details on function theory over $\mathbb{B}_{d}$, see [Rud80] and [Zhu05].
Kosinski and Vavitsas [KV23] used the Łojasiewicz inequality and some more tools from semi-analytic geometry to obtain the following classification of shift-cyclic polynomials for $d=2$. Let $\mathbb{S}_{d}$ be the boundary of $\mathbb{B}_{d}$ for $d\in\mathbb{N}$, i.e.,
@@
Here, the concept of Riesz $t$-capacity is also relevant. In fact, Vavitsas [Vav23, Theorem 16] shows that the capacity of the boundary zero set of a shift-cyclic polynomial on $\mathbb{B}_{d}$ must be 00. In a recent preprint, Vavitsas and Zarvalis [VZ24] use similar ideas and obtain the following result.
@@
Let $P\in\mathcal{P}_{d}$ be non-vanishing on $\mathbb{B}_{d}$ and suppose $\mathcal{Z}_{\mathbb{S}_{d}}(P)$ contains a real submanifold (of $\mathbb{R}^{2d}$) of real dimension $k$ with $11\leq k\leq d-$ 1, but no submanifold of higher dimension. Then, $P$ is not shift-cyclic for $t>\frac{2d-(k-1)}{2}$.
The authors point out that if $\mathcal{Z}_{\mathbb{S}_{d}}(P)$ is a finite set (in other words, $k=0$ in the statement of the above theorem), then one can argue as in the proof of Theorem 3.5 to determine its shift-cyclicity in each $\mathcal{D}_{t}(\mathbb{B}_{d})$. Thus, they conjecture that the following question has a positive answer, which classifies all shift-cyclic polynomials in the Dirichlet-type spaces over $\mathbb{B}_{d}$.
@@
Let $P\in\mathcal{P}_{d}$ be as in the hypothesis of Theorem 3.6. Then, is it true that $P$ is shift-cyclic if and only if $t\leq\frac{2d-(k-1)}{2}$?
It is worth pointing out that a result like Theorem 3.6 is not known to hold for $\mathcal{D}_{t}(\mathbb{D}^{d})$ in general, and is therefore an open area of investigation.
For the remaining discussion, we shift gears and turn to an operator theoretic approach of determining shift-cyclicity of functions. As a result, we are able to say more about multiplier-cyclicity for function spaces that may not necessarily be spaces of analytic functions (or function spaces over domains in any $\mathbb{C}^{d}$). First, we need to set up some important notation.
@@ 4.1. Reproducing kernel Hilbert spaces
Let $X$ be a set and let $\mathfrak{F}(X)$ be the space of all complex-valued functions on $X$. A space $\mathcal{H}\subset\mathfrak{F}(X)$ is said to be a reproducing kernel Hilbert space (RKHS) if it is a Hilbert space (with inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$) such that the point-evaluations
$\Lambda_{x}:f\mapsto f(x)$
are bounded linear functionals on $\mathcal{H}$. By the Riesz representation theorem, it then follows that there exist $K_{x}\in\mathcal{H}$ for each $x\in X$ such that
@@
is called the reproducing kernel (or, simply, kernel) of $\mathcal{H}$, and the functions $K_{x}$ are called the kernel functions for each $x\in X$. Using a standard orthogonality/separation argument, it is not difficult to show that
$\mathcal{H}=\mathcal{H}(K):=\overline{\operatorname{span}}\{K_{x}:x\in X\},$
and therefore, functions in a RKHS are determined by its kernel functions. Indeed, if $f\in\mathcal{H}\ominus\mathcal{H}(K)$, then $f\equiv 0$ because
@@
[AM02, Proposition 2.18] shows that the reproducing kernel $K$ of a RKHS $\mathcal{H}$ can be obtained from any orthonormal basis $\{f_{\xi}:\xi\in\Xi\}$ via
@@
An important property of every reproducing kernel is that it is a positive semi-definite (PSD) function, i.e, given any $N\in\mathbb{N}$ and any $\{x_{1},\dots,x_{N}\}\subset X$, the matrix $\{K(x_{j},x_{k})\}_{N\times N}$ is positive semi-definite. Just like matrices/operators, we write $K\geq 0$ to mean $K$ is PSD. We also note that $K$ is self adjoint, i.e.,
@@
Here, the overline denotes complex conjugate. We end this subsection with the following theorem of Moore (communicated by Aronszajn in [Aro50, Part I Section 2 (4)]), which shows that these two properties determine reproducing kernels [[, also see]Theorem 2.23]AM02.
@@
If $K:X\times X\to\mathbb{C}$ is a self adjoint PSD map, then there exists a RKHS $\mathcal{H}(K)\subset\mathfrak{F}(X)$ such that $K$ is the reproducing kernel of $\mathcal{H}(K)$.
@@ 4.1.1. Multiplier algebras
():=$\mathcal{M}(\mathcal{H}):=\{\varphi\in\mathfrak{F}(X):\varphi f\in\mathcal{H}\text{ for all }f\in\mathcal{H}\}.$
@@
This shows that every multiplier is a bounded map on $X$ [^11] We know from Section 3.1.1 that the converse is not true for the Dirichlet space.. Now, recall that if $T$ is a bounded operator on any Hilbert space, then
@@
A quick corollary of this fact is that $\varphi$ is a multiplier of $\mathcal{H}$ with $\|\varphi\|_{\mathcal{M}(\mathcal{H})}\leq c$ if and only if
@@
We wish to study the properties of functions in a RKHS through its multipliers, since they are ‘easier’ to handle in some sense. For instance, the multiplier algebra of $H^{2}(\mathbb{D})$ is $H^{\infty}(\mathbb{D})$, and we saw that the zero sets of every $H^{2}(\mathbb{D})$ is the same as that of a Blaschke product, which is an $H^{\infty}(\mathbb{D})$ function. This is also true for $\mathcal{D}$ and its multiplier algebra [[, see]Theorem 2 (c)]RS88, but not true for $L^{2}_{a}(\mathbb{D})$ [[, see]Theorem 1]Hor74, where it is easy to verify that
$\mathcal{M}(L^{2}_{a}(\mathbb{D}))=H^{\infty}(\mathbb{D}).$
The problem of determining zero sets is often related to determining interpolating sequences inside $X$. This brings us to the following question, which was first posed and answered by Pick [Pic15] (and independently by Nevanlinna [Nev19]) for $H^{\infty}(\mathbb{D})$. Given $N$ points $\{x_{1},\dots,x_{N}\}\subset X$ and $N$ target values $\{c_{1},\dots,c_{N}\}\subset\mathbb{C}$, when can we find a multiplier $\varphi\in\mathcal{M}(\mathcal{H})$ such that
@@
For $H^{2}(\mathbb{D})$, Pick [Pic15] showed that this also a sufficient condition for finite interpolating sequences. In this case, we say that $K$ has the Pick property.
The point of the above discussion is that one can construct $H^{\infty}(\mathbb{D})$ functions by just assigning target values on a finite set of points that help satisfy a certain PSD condition with the kernel of $H^{2}(\mathbb{D})$. It is now natural to ask when and for what kinds of kernels is the Pick property also sufficient.
@@ 4.1.2. Complete Pick property
For $m,n\in\mathbb{N}$, let $\mathfrak{F}(X)\otimes\mathbb{C}^{m\times n}$ be the collection of $m\times n$ matrix-valued functions on $X$ and define the multiplier algebra
$\mathcal{M}(\mathcal{H}\otimes\mathbb{C}^{n},\mathcal{H}\otimes\mathbb{C}^{m}):=\{\Phi\in\mathfrak{F}(X)\otimes\mathbb{C}^{m\times n}:\Phi F\in\mathcal{H}\otimes\mathbb{C}^{m}\text{ if }F\in\mathcal{H}\otimes\mathbb{C}^{n}\}.$
@@
where $I_{m}$ is the $m\times m$ identity matrix. If we were to then consider a matrix-valued interpolation problem, i.e., given $N$ points $\{x_{1},\dots,x_{N}\}\subset X$ and $N$ target matrices $\{C_{1},\dots C_{N}\}\subset\mathbb{C}^{m\times n}$, we wish to find some $\Phi\in\mathcal{M}(\mathcal{H}\otimes\mathbb{C}^{n},\mathcal{H}\otimes\mathbb{C}^{m})$ such that
@@
then it is clear that the following is a necessary condition for such a $\Phi$ to exist
@@
When this condition is also sufficient, we say that the kernel $K$ has the complete Pick property, and we refer to it as a complete Pick kernel.
It turns out that if $K$ satisfies some minimal conditions, then it is possible to determine if it is a complete Pick kernel. We say that $K$ is irreducible if $K(x,y)\geq 0$ for all $x,y\in X$, and $K_{x}$ and $K_{y}$ are linearly independent for any distinct $x,y\in X$ Any irreducible kernel $K$ can be normalized at a given point $x_{0}\in X$, i.e., we can replace $K$ with another irreducible kernel $\widetilde{K}$ such that $\mathcal{H}(K)$ and $\mathcal{H}(\widetilde{K})$ are unitarily equivalent and $\widetilde{K}_{x_{0}}\equiv 1$ on $X$ [[, see]Section 2.6]AM02. The following theorem of Agler and McCarthy [AM00, Theorem 4.2] gives the classification we need [[, see also]]McC92, McC94, Qui93.
@@
Let $K$ be an irreducible kernel on a set $X$ that is normalized at some $x_{0}\in X$. Then, $K$ is a complete Pick kernel if and only if there exists a Hilbert space $\mathcal{K}$ and a map $\mathfrak{b}:X\to\mathcal{K}$ such that $\|\mathfrak{b}(x)\|_{\mathcal{K}}\leq 1$ for all $x\in X$, $\mathfrak{b}(x_{0})=0$ and
@@ 4.1.3. Examples
    In order to determine whether $\mathcal{D}$ is a complete Pick space we require the following lemma, which follows from a classical result due to Kaluza [Kal28] on the coefficients of the inverse of a power-series in one variable.
@@
    The most important example of a complete Pick space is the Drury–Arveson space $\mathcal{H}^{2}_{d}$ on $\mathbb{B}_{d}$ for $d\in\mathbb{N}$. We introduced $\mathcal{H}^{2}_{d}$ in Section 3.3 as the space $\mathcal{D}_{d-1}(\mathbb{B}_{d})$. It can also be defined as a RKHS with kernel [^12] The two topologies here will be equivalent but not the same. We omit the technical details.
@@
    The Drury–Arveson space is notable because the space $\mathcal{K}$ in Theorem 4.3 can always be taken to be an infinite variable version of the Drury–Arveson space – in the same vein as $H^{2}(\mathbb{D}^{\infty}_{2})$ for the Hardy space. We shall not be needing this fact here, but the curious reader may find more details on this topic in [AM02, Chapter 8].
@@
    It is not at all easy to verify using Theorem 4.3 that $\mathcal{W}^{1,2}$ is a complete Pick space. See [AM02, Theorem 7.43] for a detailed proof of this fact.

### v3

@@ 3.3. On the unit ball $\mathbb{B}_{d}$
In this subsection, we note that the Łojasiewicz inequality and the notion of Riesz $t$-capacity helps generalize this result to a similar family of spaces over $\mathbb{B}_{2}$, and also allows us to go beyond the $d=2$ case to obtain some partial results on shift-cyclicity of polynomials.
Let $d\in\mathbb{N}$ be fixed. For each $\alpha\in\mathbb{Z}_{+}^{d}$, we write
@@
Now, for each $t\in\mathbb{R}$, we define the Dirichlet-type spaces on $\mathbb{B}_{d}$ as
@@
[…]
Kosinski and Vavitsas [64] used the Łojasiewicz inequality and some more tools from semi-analytic geometry to obtain the following classification of shift-cyclic polynomials for $d=2$. Let $\mathbb{S}_{d}$ be the boundary of $\mathbb{B}_{d}$ for $d\in\mathbb{N}$, i.e.,
@@
Here, the concept of Riesz $t$-capacity is also relevant. In fact, Vavitsas [104, Theorem 16] shows that the capacity of the boundary zero set of a shift-cyclic polynomial on $\mathbb{B}_{d}$ must be 00. In a recent preprint, Vavitsas and Zarvalis [105] use similar ideas and obtain the following result.
@@
Let $P\in\mathcal{P}_{d}$ be non-vanishing on $\mathbb{B}_{d}$ and suppose $\mathcal{Z}_{\mathbb{S}_{d}}(P)$ contains a real submanifold (of $\mathbb{R}^{2d}$) of dimension $k$ with $11\leq k\leq d-$ 1, but no submanifold of higher dimension. Then, $P$ is not shift-cyclic in $\mathcal{D}_{t}(\mathbb{B}_{d})$ for $t>\frac{2d-(k-1)}{2}$.
The authors point out that if $\mathcal{Z}_{\mathbb{S}_{d}}(P)$ is a finite set (in other words, $k=0$ in the statement of the above theorem), then one can argue as in the proof of Theorem 3.5 to determine its shift-cyclicity in each $\mathcal{D}_{t}(\mathbb{B}_{d})$. Thus, they conjecture that the following question has a positive answer, which then classifies all shift-cyclic polynomials in the Dirichlet-type spaces over $\mathbb{B}_{d}$.
@@
Let $P\in\mathcal{P}_{d}$ be as in the hypothesis of Theorem 3.6. Then, is it true that $P$ is shift-cyclic in $\mathcal{D}_{t}(\mathbb{B}_{d})$ if and only if $t\leq\frac{2d-(k-1)}{2}$?
@@ 4. Complete Pick spaces
For the remaining discussion, we shall shift gears and turn to an operator theoretic approach of studying shift-cyclicity of functions. As a result, we are able to say more about multiplier-cyclicity for function spaces that may not necessarily be spaces of analytic functions (or function spaces over domains in $\mathbb{C}^{d}$ for any $d\in\mathbb{N}$). First, we need to set up some important notation.
@@ 4.1. Reproducing kernel Hilbert spaces
Let $X$ be a set and let $\mathcal{F}(X,\mathbb{C})$ be the space of all complex-valued functions on $X$. A space $\mathcal{H}\subset\mathcal{F}(X,\mathbb{C})$ is said to be a reproducing kernel Hilbert space (RKHS) if it is a Hilbert space (with inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$) such that the point-evaluations $\Lambda_{x}:f\mapsto f(x)$ are bounded linear functionals on $\mathcal{H}$. By the Riesz representation theorem, it then follows that there exist $K_{x}\in\mathcal{H}$ for each $x\in X$ such that
@@
is called the reproducing kernel (or, simply, kernel) of $\mathcal{H}$, and the functions $K_{x}$ are called the kernel functions for each $x\in X$. Using a standard orthogonality/separation argument, it is not difficult to show that $\mathcal{H}=\mathcal{H}(\mathcal{K})$, where
$\mathcal{H}(K):=\overline{\operatorname{span}}\{K_{x}:x\in X\},$
Indeed, if $f\in\mathcal{H}\ominus\mathcal{H}(K)$ then $f\equiv 0$, because
@@
[2, Proposition 2.18] shows that the reproducing kernel $K$ of a RKHS $\mathcal{H}$ can be obtained from any orthonormal basis $\{f_{\xi}:\xi\in\Xi\}$ via
@@
An important property of every reproducing kernel is that it is a positive semi-definite (PSD) function, i.e, given any $N\in\mathbb{N}$ and any $\{x_{1},\dots,x_{N}\}\subset X$, the matrix $(K(x_{j},x_{k}))_{N\times N}$ is PSD. Just like matrices/operators, we write $K\geq 0$ to mean $K$ is PSD. We also note that $K$ is self adjoint, i.e.,
@@
Here, the overline denotes complex conjugate. We end this subsection with the following theorem of Moore (communicated by Aronszajn in [6, Part I Section 2 (4)]), which shows that these two properties determine reproducing kernels (see also [2, Theorem 2.23]).
@@
If $K:X\times X\to\mathbb{C}$ is a self adjoint PSD map, then there exists a RKHS $\mathcal{H}(K)\subset\mathcal{F}(X,\mathbb{C})$ such that $K$ is the reproducing kernel of $\mathcal{H}(K)$.
@@ 4.1.1. Multiplier algebras
():=$\mathcal{M}(\mathcal{H}):=\{\varphi:X\to\mathbb{C}:\varphi f\in\mathcal{H}\text{ for all }f\in\mathcal{H}\}.$
@@
This shows that every multiplier is a bounded function on $X$ [^12] We know from Section 3.1.1 that the converse is not true for the Dirichlet space.. Now, recall that if $T$ is a bounded operator on any Hilbert space, then
@@
A quick corollary of this fact is that $\varphi\in\mathcal{M}(\mathcal{H})$ with $\|\varphi\|_{\mathcal{M}(\mathcal{H})}\leq c$ if and only if
@@
We wish to study the properties of functions in a RKHS through its multipliers. For instance, the multiplier algebra of $H^{2}(\mathbb{D})$ is $H^{\infty}(\mathbb{D})$, and we saw that the zero sets of every $H^{2}(\mathbb{D})$ is the same as that of a Blaschke product, which is an $H^{\infty}(\mathbb{D})$ function. This is also true for $\mathcal{D}$ and its multiplier algebra (see [89, Theorem 2 (c)]), but not true for $L^{2}_{a}(\mathbb{D})$, where $\mathcal{M}(L^{2}_{a}(\mathbb{D}))=H^{\infty}(\mathbb{D})$ (see [54, Theorem 1]). The problem of determining zero sets is often related to determining interpolating sequences. This brings us to the following question, which was first posed and answered by Pick [86] (and independently by Nevanlinna [80]) for $H^{\infty}(\mathbb{D})$. Given $N$ points $\{x_{1},\dots,x_{N}\}\subset X$, $N$ target values $\{c_{1},\dots,c_{N}\}\subset\mathbb{C}$ and some $c>0$, when can we find a multiplier $\varphi\in\mathcal{M}(\mathcal{H})$ with $\|\varphi\|_{\mathcal{M}(\mathcal{H})}\leq c$ such that
@@
For $H^{2}(\mathbb{D})$, Pick [86] showed that this condition is also sufficient for the existence of such a $\varphi$. In this case, we say that $K$ has the Pick property.
The point of the above discussion is that one can construct $H^{\infty}(\mathbb{D})$ functions with a norm bound by just assigning target values on a finite set of points that help satisfy a certain PSD condition with the kernel of $H^{2}(\mathbb{D})$. It is now natural to ask when and for what kinds of kernels is the Pick property also sufficient.
@@ 4.1.2. Complete Pick property
For $m,n\in\mathbb{N}$, let $\mathcal{F}(X,\mathbb{C})\otimes\mathbb{C}^{m\times n}$ be the collection of $m\times n$ matrix-valued functions on $X$ and define the multiplier algebra
$\mathcal{M}(\mathcal{H}\otimes\mathbb{C}^{n},\mathcal{H}\otimes\mathbb{C}^{m}):=\{\Phi\in\mathcal{F}(X,\mathbb{C})\otimes\mathbb{C}^{m\times n}:\Phi F\in\mathcal{H}\otimes\mathbb{C}^{m}\text{ if }F\in\mathcal{H}\otimes\mathbb{C}^{n}\}.$
@@
where $I_{m}$ is the $m\times m$ identity matrix. Let us then consider a matrix-valued interpolation problem, i.e., given $N$ points $\{x_{1},\dots,x_{N}\}\subset X$, $N$ target matrices $\{C_{1},\dots C_{N}\}\subset\mathbb{C}^{m\times n}$ and some $C>0$, we wish to find a multiplier $\Phi\in\mathcal{M}(\mathcal{H}\otimes\mathbb{C}^{n},\mathcal{H}\otimes\mathbb{C}^{m})$ with norm at most $C$ such that
@@
It is clear that the following is a necessary condition for such a $\Phi$ to exist
@@
When this condition is also sufficient for the existence of such a $\Phi$ with norm at most $C$, we say that the kernel $K$ has the complete Pick property, and we refer to it as a complete Pick kernel.
It turns out that if $K$ satisfies some minimal conditions, then it is possible to determine if it is a complete Pick kernel. We say that $K$ is irreducible if $K(x,y)\neq 0$ for all $x,y\in X$, and $K_{x}$ and $K_{y}$ are linearly independent for any distinct $x,y\in X$ Any irreducible kernel $K$ can be normalized at a given point $x_{0}\in X$, i.e., we can replace $K$ with another irreducible kernel $\widetilde{K}$ such that $\mathcal{H}(K)$ and $\mathcal{H}(\widetilde{K})$ are unitarily equivalent and $\widetilde{K}_{x_{0}}\equiv 1$ on $X$ (see [2, Section 2.6]). The following theorem of Agler and McCarthy [1, Theorem 4.2] gives the classification we need (also see [74, 75, 87]).
@@
Let $K$ be an irreducible kernel on a set $X$ that is normalized at some $x_{0}\in X$. Then, $K$ is a complete Pick kernel if and only if there exists a Hilbert space $\mathcal{K}$ and a map $\mathfrak{b}:X\to\mathcal{K}$ such that $\|\mathfrak{b}(x)\|_{\mathcal{K}}<1$ for all $x\in X$, $\mathfrak{b}(x_{0})=0$ and
@@ 4.1.3. Examples
    In order to determine whether $\mathcal{D}$ is a complete Pick space we require the following lemma, which follows from a classical result due to Kaluza [59] on the coefficients of the inverse of a power-series in one variable.
@@
    The most important example of a complete Pick space is the Drury–Arveson space $\mathcal{H}^{2}_{d}$ on $\mathbb{B}_{d}$ for $d\in\mathbb{N}$. We introduced $\mathcal{H}^{2}_{d}$ in Section 3.3 as the space $\mathcal{D}_{d-1}(\mathbb{B}_{d})$. It can also be defined as the RKHS with kernel [^13] Note that both these definitions result in different inner products, but it can be checked that they generate the same topology. We omit the technical details.
@@
    An infinite variable version of the Drury–Arveson space is notable for being the ‘universal’ complete Pick space. In essence, if $\mathcal{H}(K)$ is separable, then the space $\mathcal{K}$ in Theorem 4.3 can instead be chosen to be $\ell^{2}(\mathbb{Z}_{+})$, and the map $\mathfrak{b}$ can be chosen to be an embedding of $X$ into $\ell^{2}(\mathbb{Z}_{+})$. If we now construct the Drury–Arveson space on the unit ball of $\ell^{2}(\mathbb{Z}_{+})$, i.e. the RKHS over the unit ball of $\ell^{2}(\mathbb{Z}_{+})$ with kernel
    $K_{\ell^{2}}(v,w):=\frac{1}{1-\langle v,w\rangle_{\ell^{2}(\mathbb{Z}_{+})}},$
    then the map
    $K_{x}\mapsto[K_{\ell^{2}}]_{\mathfrak{b}(x)}$
    gives an isometric linear embedding of $\mathcal{H}(K)$ into $\mathcal{H}(K_{\ell^{2}})$. One may think of $\mathcal{H}(K_{\ell^{2}})$ as an infinite variable version of the Drury–Arveson space. The reader may find more details on this topic in [2, Chapter 8].
    It is also worth noting that infinitely many variables are needed in general, since the Dirichlet space, for instance, cannot be embedded in $H^{2}_{d}$ for any $d\in\mathbb{N}$ (see [48, Corollary 11.9]). Moreover, Hartz [49] later showed that this cannot be achieved even under a weaker hypothesis.
@@
    It is not at all easy to verify using Theorem 4.3 that $\mathcal{W}^{1,2}$ is a complete Pick space. See [2, Theorem 7.43] for a detailed proof of this fact.